The fuel for a gas snowblower consists of 1.3 parts oil to 24.7 parts gas. Determine the ratio of oil to gas in simplest form.
Answer by dabanfield(803)

The fuel for a gas snowblower consists of 1.3 parts oil to 24.7 parts gas. Determine the ratio of oil to gas in simplest form.
Ratio = 1.3/24.7 = 13/247 = 0.052631579